# A Novel Approach to Reduce Derivative Costs in Variational Quantum Algorithms

### v1

@@ Abstract
We present a detailed numerical study of an alternative approach, named Quantum Non-Demolition Measurement (QNDM) Solinas et al. (2023), to efficiently estimate the gradients or the Hessians of a quantum observable. This is a key step and a resource-demanding task when we want to minimize the cost function associated with a quantum observable. In our detailed analysis, we account for all the resources needed to implement the QNDM approach with a fixed accuracy and compare them to the current state-of-the-art method Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We find that the QNDM approach is more efficient, i.e. it needs fewer resources, in evaluating the derivatives of a cost function.These advantages are already clear in small dimensional systems and are likely to increase for practical implementations and more realistic situations. Since the vast majority of the Variational Quantum Algorithms can be formulated in the discussed framework, our results can have significant implications in quantum optimization algorithms and make the QNDM approach a valuable alternative to implement Variational Quantum Algorithms on near-term quantum computers.
@@ I Introduction
The classical optimization procedures are well-known and established Boyd and Vandenberghe (2004); Adby (2013) and they usually exploit information on the derivatives (gradient) of the cost function to update the parameters. At the quantum level, the way to extract the desired information from a quantum system is still debated and open Banchi and Crooks (2021); Wierichs et al. (2022). Indeed, the derivative cannot be directly associated with a Hermitian operator Solinas and Gasparinetti (2015, 2016) and, therefore, there is no unique way to measure it. Still, this is a crucial step since it is usually a bottleneck of the computation with quantum computers. The most straightforward and used approach to obtain the gradient of the cost function is to run a quantum circuit and measure a quantum observable at a certain point in the parameter space. Then, we repeat the procedure changing the circuit parameters by a little along each direction, separately Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We call this method, the Direct Measurements (DM) approach and, for every point in the parameter space, it requires two observable measurements to determine each component of the gradient.
In this article, we analyse a novel method to compute the derivatives of the cost function implemented on the quantum circuit. This is called the Quantum Non-Demolition Measurement (QNDM) approach and it was first theoretically described in Ref. Solinas et al. (2023). The main idea is to extract the information about the derivative with a single measurement of a quantum observable. This is obtained by coupling a quantum detector to the original quantum system so that the information about the value of the derivative of the cost function is stored in the phase of the detector. Finally, the phase is measured to extract this information Solinas et al. (2023).
@@
The paper is structured as follows. In Sec. II, we recall the basic ideas of VQA and the two approaches to derivative evaluation. Section III is devoted to the analysis of the statistical errors and the resource cost. Section IV.2 presents the numerical comparison, while Sec.VI contains the conclusions. The codes to calculate derivatives with QNDM used within this article is accessible via GitHub S.Caletti and G.Minuto (2024). In Appendix B we show how to install the QNDM package.
@@ II.1 General Settings
In many applications McArdle et al. (2020); Tilly et al. (2022), the observable $\hat{M}$ can be written as a weighted sum of Pauli strings $\hat{P}_{i}=\prod_{j}\hat{A}_{i}^{j}$, where $\hat{A}_{j}\in[X_{j},Y_{j},Z_{j}]$ is one of the usual Pauli operators acting on the $j$-th qubit. With this notation, we have that
@@ II.3 Quantum Non-Demolition Measurement
where ${\rm Tr}_{S}$ denotes the trace over the system degrees of freedom. By direct comparison, it is evident that (11) is proportional to Eq. (8). Thus, both the QNDM and the DM approach can be used to compute the derivative of $f(\vec{\theta})$.
@@ III.1 Cost analysis and comparison between DM and QNDM approaches
This quadratic behavior in $J$ arises from two components: firstly, the inherent linear advantage that QNDM  holds over DM  Solinas et al. (2023), and secondly, the dependency of the MSEDM on the number of $J$, which is absent in MSEQNDM.
This regime is particularly relevant for quantum chemistry simulations. For instance, in simulations of moderately complex molecules Wecker et al. (2015), we have $k\approx 10^{9}-10^{10}$ operations to implement $U(\vec{\theta})$, $n\approx 10^{2}-10^{3}$ (the expected dimension of quantum computers in the next years), and $J>10^{3}$ as the number of Pauli strings in the Hamiltonian. With these numbers, $k\gg nJ$, resulting in a reduction of roughly $J^{2}$ resources when using the QNDM approach. This advantage is likely to increase for larger and more complex molecules.
@@ IV.1 Errors Analysis
Figure 4: The main plot represents the value of the derivative over the $j$-th direction $\mu(g_{j}^{i})$ [$i=$DM (red markers) and $i=$QNDM (blue markers)], averaged over a collection of different realizations, and shown as a function of the number of Pauli strings. In inset we show the ratio between the $MSE_{DM}/MSE_{QNDM}$. The simulated quantum system has $n=10$ qubits and the number of shots is fixed at $N=10^{4}$. Each point is averaged for $L=100$ realizations implemented as discussed in the main text.
In Fig. 4 we plot the averaged derivative $\mu(g_{j}^{i})$ over a randomly chosen direction $j$, and their errors as a function of the number of Pauli strings $J$. The DM  and QNDM  results corresponds to the red and blue markers, respectively. In this case, the average of $\mu(g_{j}^{i})$ is performed over a collection of different realizations averaged over the above parameter (i) to (iv). For each point, we randomly selected $L=100$ realizations. Also, the errors associated with the points are the result of the average of the errors of the corresponding realizations.
@@ IV.2 Costs Comparison
For practical situations, it is of paramount importance to quantitatively estimate any advantage of the QNDM algorithm over DMand in which conditions this is reached. From the discussion in Sec. III.1, we expect the QNDM approach to perform better, i.e., that it needs fewer resources, in some specific but interesting regimes. Aiming in this direction, we present a full numerical analysis of the resources needed to calculate the derivative of the cost function.
@@
Figure 5: Regime $k\gg nJ$. $a)$ The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculate for each realization by (27), such that the MSE errors are equal for both methods.
@@
Figure 6: Regime $k\ll nJ$. The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$ ($i=QNDM$(orange line) and $i=DM$(blue line)). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the product $nJ$ for a fixed number of logical operators $k=0.5\times 10^{3}$. $b)$ Ratio between DM and QNDM resources $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of the number of logical operators $k$ each colored line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculate for each realization by (27), such that the MSE errors are equal for both methods.
The regime in which $k\ll nJ$ is analyzed in Fig. 6. Panel a), on the left, shows the resource cost function $\mathcal{C}(g_{j}^{i})$ ($i=QNDM,DM$) obtained to estimate the derivative along the direction $j$. The values are plotted as a function $nJ$. The simulations are done for $n=10$ qubits with $k=0.5\cdot 10^{3}$ fixed. Each value in Fig. 6 is calculated by performing an average over a collection of different realizations, as discussed above. As delineated in tab. 1, for DM (the blu line) the number of resources increases quadratically in the number of $J$. Already for $kJ\sim 10^{3}$, the DM method need $10^{4}$ more resources than QNDM one.
@@
Notice that in all our simulations and both regimes, we have considered relatively small and simple systems to give a more conservative estimate. We might expect a further increasing advantage in realistic computational problems since they usually require more logical gates, and we deal with more complex observable $\hat{M}$.
@@ VI Conclusions
In this paper, we have presented a detailed study of an alternative approach to evaluate the derivatives of a cost function with a quantum computer. This method is called Quantum Non-Demolition Measurement (QNDM) and was first presented in Ref.Solinas et al. (2023). In addition to a quantitative estimate of the resources needed to run the QNDM protocol, we have proposed an in-depth comparison with the state-of-the-art method, to which we refer as the Direct Measurement (DM) approach Cerezo et al. (2021).
The results obtained are encouraging and pave the way for the first application of the QNDM algorithm to real applications. The QNDM approach exhibits advantages concerning the DM one in terms of resources. This advantage can be traced back to two main features. First, the use of a quantum detector to store information about the derivative of the cost function allows us to run the quantum circuit and measure only once. On the contrary, the DM approach needs two circuits and two measures (four circuits and four measures in the Hessian case).
Second, for the QNDM , the statistical uncertainty, i.e., the mean square error, results independent of a number of Pauli strings $J$ appearing in the quantum observable while, for the DM, it increases linearly with $J$. As a consequence, to have the same accuracy in the average values of the cost function, for quantum observables composed by numerous Pauli strings $J$, the number of quantum circuit runs in the DM  must increase.
The numerical simulations performed through the Qiskit framework, have confirmed the analytical estimates. Already for small systems and simplified quantum circuits, we observe a considerable reduction of the resources needed to run the QNDM  approach with respect to the DM one. We expect that this advantage further increases for more practical and interesting problems, such as the simulation of chemical compounds or molecules or small-size quantum systems Weber et al. (2014).
These results make the QNDM approach a valuable alternative to implement VQAs in the next generation of noisy quantum computers.
The Qiskit code used for the simulation is publicly available via GitHub S.Caletti and G.Minuto (2024).
@@ Acknowledgements.
SC would like to thank the University of Amsterdam and the Delta-ITP programme for support during the initial part of this work.
@@ A.1 Errors Analysis for Second Derivative
The MSEs for the second derivatives for both the methods, respectively, are equal to
@@
Notice that the differences with the first derivative MSEs in Eqs. (26) and (22) are only in the denominator and that for the second derivative, the MSEDM grows linearly with the number of Pauli strings $J$. Consequently, to achieve the same precision level as MSEQNDM in the DM approach, we need to increase the number of shots as a function of $J$. More specifically, we have that
@@
Where, similarly to to Eq. (27), we keep the linearity in $J$, only with a modified multiplicative factor due to the differences in the denominators between the MSE for the first and the second derivatives.
@@ A.2 Cost Simulations for Second Derivative
In Section V.1, we found the asymptotic cost ratios $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ for the two regimes: $k\ll nJ$ Eq. (31) and $k\gg nJ$ Eq. (30). In Figs. S4 and S3, we include simulations that illustrate how the theoretical results are also maintained for the second derivatives. For each value presented in the figures we averaged over $L=50$ different realizations.
Figure S3: Regime $k\gg nJ$. $a)$ Number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of logical operators $k$ for a fixed $nJ=300$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculate for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S3 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in function of the logical operators $k$, when we consider the limit of $k\gg nJ$. The simulations are done for $n=10$ qubits, fixed $nJ=300$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of the number of Pauli strings $J$ and for $k$ different numbers of logical operators in $U(\vec{\theta})$.
Figure S4: $a)$ The averaged number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of Pauli strings $J$ for a fixed number of logical operators $k=0.5\cdot 10^{3}$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of logical operators $k$ each coloured line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculate for each realization by (S7), such that the MSE errors are equal for both methods.

### v2

@@ Abstract
We present a detailed numerical study of an alternative approach, named Quantum Non-Demolition Measurement (QNDM) Solinas et al. (2023), to efficiently estimate the gradients or the Hessians of a quantum observable. This is a key step and a resource-demanding task when we want to minimize the cost function associated with a quantum observable. The statistical study of the error leads to a further performance improvement with respect to the original proposal in Ref. Solinas et al. (2023). In our detailed analysis, we account for all the resources needed to implement the QNDM approach with a fixed accuracy and compare them to the current state-of-the-art method Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We find that the QNDM approach is more efficient, i.e. it needs fewer resources, in evaluating the derivatives of a cost function. These advantages are already clear in small dimensional systems and are likely to increase for practical implementations and more realistic situations. A significant outcome of our study is the implementation of the QNDM method in Python, provided in the supplementary material S.Caletti and G.Minuto (2024). Given that most Variational Quantum Algorithms can be formulated within this framework, our results can have significant implications in quantum optimization algorithms and make the QNDM approach a valuable alternative to implement Variational Quantum Algorithms on near-term quantum computers.
@@ I Introduction
The classical optimization procedures are well-known and established Boyd and Vandenberghe (2004); Adby (2013) and they usually exploit information on the derivatives (gradient) of the cost function to update the parameters. At the quantum level, the way to extract the desired information from a quantum system is still debated and open Banchi and Crooks (2021); Wierichs et al. (2022). Indeed, the derivative cannot be directly associated with a Hermitian operator Solinas and Gasparinetti (2015, 2016) and, therefore, there is no unique way to measure it. Still, this is a crucial step since it usually causes a bottleneck in computation with quantum computers. The most straightforward and used approach to obtain the gradient of the cost function is to run a quantum circuit and measure a quantum observable at a certain point in the parameter space. Then, we repeat the procedure changing the circuit parameters by a little along each direction, separately Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We call this method, the Direct Measurements (DM) approach and, for every point in the parameter space, it requires two observable measurements to determine each component of the gradient.
In this article, we analyze a novel method to compute the derivatives of the cost function implemented on the quantum circuit. This is called the Quantum Non-Demolition Measurement (QNDM) approach and it was first theoretically described in Ref. Solinas et al. (2023). The main idea is to extract the information about the derivative with a single measurement of a quantum observable. This is obtained by coupling a quantum detector to the original quantum system so that the information about the value of the derivative of the cost function is stored in the phase of the detector. Finally, the phase is measured to extract this information Solinas et al. (2023).
@@
The paper is structured as follows. In Sec. II, we recall the basic ideas of VQA and the two approaches to derivative evaluation. Section III is devoted to the analysis of the statistical errors and the resource cost. Section IV.2 presents the numerical comparison, while Sec.VI contains the conclusions. The codes to calculate derivatives with QNDM used within this article are accessible via GitHub S.Caletti and G.Minuto (2024). In Appendix B we show how to install the QNDM package.
@@ II.1 General Settings
In many applications McArdle et al. (2020); Tilly et al. (2022), the observable $\hat{M}$ can be written as a weighted sum of Pauli strings $\hat{P}_{i}=\prod_{j}\hat{A}_{i}^{j}$, where $\hat{A}_{j}\in[X_{j},Y_{j},Z_{j},I_{j}]$ is one of the usual Pauli operators acting on the $j$-th qubit. With this notation, we have that
@@ II.3 Quantum Non-Demolition Measurement
where ${\rm Tr}_{S}$ denotes the trace over the system degrees of freedom. By direct comparison, Eq. (11) is proportional to Eq. (8),
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}=2\sin(s)g_{j}.$ (12)
Thus, both the QNDM and the DM approach can be used to compute the derivative of $f(\vec{\theta})$.
@@ III.1 Cost analysis and comparison between DM and QNDM approaches
This behavior shows that the cost to calculate a derivative with DM is $J^{2}$ higher than QNDM so that, in this regime, we have quadratic speed up with the number of Pauli strings. This arises from two components: firstly, the inherent linear advantage that QNDM  holds over DM  Solinas et al. (2023), and secondly, the dependency of the MSEDM on the number of $J$, which is absent in MSEQNDM. This regime is particularly relevant for quantum chemistry simulations. For instance, in simulations of moderately complex molecules Wecker et al. (2015), we have $k\approx 10^{9}-10^{10}$ operations to implement $U(\vec{\theta})$, $n\approx 10^{2}-10^{3}$ (the expected dimension of quantum computers in the next years), and $J>10^{3}$ as the number of Pauli strings in the Hamiltonian. With these numbers, $k\gg nJ$, resulting in a reduction of roughly $J^{2}$ resources when using the QNDM approach. This advantage is likely to increase for larger and more complex molecules.
@@ IV.1 Errors Analysis
Figure 4: The main plot represents the value of the derivative over the $j$-th direction $\mu(g_{j}^{i})$ [$i=$DM (red markers) and $i=$QNDM (blue markers)], averaged over a collection of different realizations, and shown as a function of the number of Pauli strings. In the inset, we show the ratio between the $MSE_{DM}/MSE_{QNDM}$. The simulated quantum system has $n=10$ qubits and the number of shots is fixed at $N=10^{4}$. Each point is averaged for $L=100$ realizations implemented as discussed in the main text.
In Fig. 4 we plot the averaged derivative $\mu(g_{j}^{i})$ over a randomly chosen direction $j$, and their errors as a function of the number of Pauli strings $J$. The DM  and QNDM  results correspond to the red and blue markers, respectively. In this case, the average of $\mu(g_{j}^{i})$ is performed over a collection of different realizations averaged over the above parameter (i) to (iv). For each point, we randomly selected $L=100$ realizations. Also, the errors associated with the points are the result of the average of the errors of the corresponding realizations.
@@ IV.2 Costs Comparison
For practical situations, it is of paramount importance to quantitatively estimate any advantage of the QNDM algorithm over DMand in which conditions this is reached. From the discussion in Sec. III.1, we expect the QNDM approach to perform better, i.e., it needs fewer resources, in some specific but interesting regimes. Aiming in this direction, we present a full numerical analysis of the resources needed to calculate the derivative of the cost function.
@@
Figure 5: Regime $k\gg nJ$. $a)$ The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
@@
Figure 6: Regime $k\ll nJ$. The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$ ($i=QNDM$(orange line) and $i=DM$(blue line)). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the product $nJ$ for a fixed number of logical operators $k=0.5\times 10^{3}$. $b)$ Ratio between DM and QNDM resources $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of the number of logical operators $k$ each colored line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
The regime in which $k\ll nJ$ is analyzed in Fig. 6. Panel a), on the left, shows the resource cost function $\mathcal{C}(g_{j}^{i})$ ($i=QNDM,DM$) obtained to estimate the derivative along the direction $j$. The values are plotted as a function $nJ$. The simulations are done for $n=10$ qubits with $k=0.5\cdot 10^{3}$ fixed. Each value in Fig. 6 is calculated by performing an average over a collection of different realizations, as discussed above. As delineated in tab. 1, for DM (the blue line) the number of resources increases quadratically in the number of $J$. Already for $kJ\sim 10^{3}$, the DM method need $10^{4}$ more resources than QNDM one.
@@
Notice that the results obtained complement the analytic analysis provided in Solinas et al. (2023). Specifically, the error analysis, Sec. III and Sec. IV.1, shows that, for the QNDM, the statistical uncertainty, i.e., the mean square error, is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM, it increases linearly with $J$. This ensures additional savings in resources, as shown in Fig. 5 and Fig. 6, compared to what was anticipated in our analytical analysis Solinas et al. (2023). Additionally, in all our simulations and both regimes, we have considered relatively small and simple systems to provide a more conservative estimate. We might expect a further increasing advantage in realistic computational problems, as they usually require more logical gates and involve more complex observables $\hat{M}$.
@@ VI Conclusions
In this paper, we have presented a detailed study of an alternative approach to evaluating the derivatives of a cost function using a quantum computer, along with its implementation in Python publicly available via GitHub S.Caletti and G.Minuto (2024). This method called Quantum Non-Demolition Measurement (QNDM), was first introduced in Ref. Solinas et al. (2023). In addition to providing a quantitative estimate of the resources needed to run the QNDM protocol, we have conducted an in-depth comparison with the state-of-the-art Direct Measurement (DM) approach Cerezo et al. (2021). The results obtained are encouraging and pave the way for the first applications of the QNDM  algorithm to real-world problems since it allows us to consistently reduce the resources needed for the computation.
This advantage can be attributed to two main features. First, as highlighted in Solinas et al. (2023), the use of a quantum detector to store information about the derivative of the cost function allows us to run the quantum circuit and measure only once. In contrast, the DM approach requires two circuits and two measurements (four circuits and four measurements for the Hessian). Second, introduced here for the first time, the statistical uncertainty (i.e., the mean square error) for the QNDM is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM approach, it increases linearly with $J$. Consequently, to achieve the same accuracy in the average values of the cost function for quantum observables composed of numerous Pauli strings $J$, the number of quantum circuit runs in the DM approach must increase.
The numerical simulations performed using the Qiskit framework have confirmed the analytical estimates. Even for small systems and simplified quantum circuits, we observe a considerable reduction in the resources needed to run the QNDM approach compared to the DM method. We anticipate that this advantage will further increase for more practical and complex problems, such as the simulation of chemical compounds, molecules, or small quantum systems Weber et al. (2014). These results position the QNDM approach as a valuable alternative for implementing Variational Quantum Algorithms (VQAs) on the next generation of noisy quantum computers.
@@ Acknowledgements.
SC would like to thank the University of Amsterdam and the Delta-ITP program for their support during the initial part of this work.
@@ A.1 Errors Analysis for Second Derivative
The MSEs for the second derivatives for both methods, respectively, are equal to
@@
Notice that the differences with the first derivative MSEs in Eqs. (27) and (23) are only in the denominator and for the second derivative, the MSEDM grows linearly with the number of Pauli strings $J$. Consequently, to achieve the same precision level as MSEQNDM in the DM approach, we need to increase the number of shots as a function of $J$. More specifically, we have that
@@
Where, similarly to Eq. (28), we keep the linearity in $J$, only with a modified multiplicative factor due to the differences in the denominators between the MSE for the first and the second derivatives.
@@ A.2 Cost Simulations for Second Derivative
In Section V.1, we found the asymptotic cost ratios $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ for the two regimes: $k\ll nJ$ Eq. (32) and $k\gg nJ$ Eq. (31). In Figs. S4 and S3, we include simulations that illustrate how the theoretical results are also maintained for the second derivatives. For each value presented in the figures, we averaged over $L=50$ different realizations.
Figure S3: Regime $k\gg nJ$. $a)$ Number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of logical operators $k$ for a fixed $nJ=300$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S3 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in the function of the logical operators $k$, when we consider the limit of $k\gg nJ$. The simulations are done for $n=10$ qubits, fixed $nJ=300$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of the number of Pauli strings $J$ and for $k$ different numbers of logical operators in $U(\vec{\theta})$.
Figure S4: $a)$ The averaged number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of Pauli strings $J$ for a fixed number of logical operators $k=0.5\cdot 10^{3}$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of logical operators $k$ each coloured line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.